# Dark Matter Density Profile Around a Newborn First Star

### v1

@@ 1 Introduction
Gravitational wave (GW) astronomy offers a new avenue to probe fundamental physics [4]. Dynamical friction exerted by ambient DM on a black‑hole (BH) binary perturbs its orbital evolution and imprints a characteristic phase shift on the emitted GW signal [5]. If the central BH is surrounded by concentrated DM, dubbed a spike [6], dress [7], or mound [8], the accumulated phase shift can be significant. State‑of‑the‑art waveform calculations suggest that next‑generation facilities are expected to have enough sensitivities to detect the effects [9, 10, 11]; e.g., third-generation ground-based (ET/CE) and space-based (LISA-like) detectors. Future GW observatories can thus serve as indirect DM detectors, but the feasibility heavily relies on the model of the DM density profile.
@@
We present the result of very high resolution cosmological hydrodynamical simulations that resolve the DM distribution inside a minihalo without performing particle splitting through linear density perturbations $z=49$ to gravitational collapse. Our flagship run (labeled L3) attains a DM particle mass of $m_{\rm dm,min}=3.66\times 10^{-4}\,M_{\odot}$, and thus improves mass resolution by a factor $\approx 130$ relative to the previous works [25, 26]. We find that initially a slowly rotating core is formed. As the baryonic cloud contracts and the local Keplerian velocity increases, the DM core undergoes further contraction, developing a steeper inner slope. We derive a semi‑analytic formula for the DM density distribution at the Pop III star formation by fitting the simulation result and by extrapolating to protostellar sizes. The result provides a realistic input for GW waveform models with the contribution of concentrated DM.
2 Methods
We generate a hierarchical zoom-in cosmological initial conditions using MUSIC [27] for a volume of $L_{\rm box}=100$ comoving kiloparsec (ckpc) per side at $z_{\rm ini}=499$. We adopt $\Lambda$-Cold Dark Matter ($\Lambda$CDM) cosmology with the cosmological parameters $\Omega_{\rm m}=0.31$, $\Omega_{\rm b}=0.048$, $\Omega_{\Lambda}=0.69$, $H_{0}=68\,{\rm km\,s^{-1}}\,{\rm Mpc}^{-1}$, and $n_{\rm s}=0.96$. To accelerate the formation of nonlinear objects (halos) in a small computational domain, we artificially increased the root-mean-square matter fluctuation averaged over a sphere of radius $8\,h^{-1}\,$pc from $\sigma_{8}=0.83$ to $\sigma_{8}=2$. This model is dubbed L0, in which the finest particle masses are $m_{\rm dm,min}=0.805\,M_{\odot}$ for DM and $m_{\rm b,min}=0.145\,M_{\odot}$ for baryon (gas) component.
We perform a set of cosmological simulations with the parallel $N$-body/smoothed particle Hydrodynamics (SPH) code GADGET-3 [28], adapted for metal-free star formation with including detailed non-equilibrium chemistry of 14 species (e-, H, H+, H-, He, He+, He++, H2, H+${}_{2}^{+}$, D, D+, HD, HD+, HD-) [29]. We adopt a two percent of the initial inter-particle distance as the gravitational softening length as $\varepsilon=L_{\rm box}/N_{\rm eff}^{1/3}/50$. We run model L0 until the baryon density reaches $\rho_{\rm b}=10^{6}\,{\rm GeV\,cm^{-3}}$, when the primordial gas cloud forms and collapses. At this stage, the DM minihalo hosting the gas cloud has a virial radius of approximately $R_{\rm virial}\simeq 100\,$pc and a mass of $M_{\rm virial}\simeq 3\times 10^{5}\,M_{\odot}$ at redshift $z_{\rm col}\simeq 22$, representing the typical properties of a minihalo for the Pop III star formation [17, 30].
@@ 2 Methods
Table 1: Column 1: model name; Column 2: mass of the finest DM ($N$-body) particle; Column 3: mass of the finest baryon/gas (SPH) particle; Column 4: gravitational softening length of the finest DM particle; and Column 5: redshift when the gas cloud collapses ($\rho_{\rm b,max}=10^{6}\,{\rm GeV\,cm^{-3}}$).
Next, we mark the DM and gas particles within the target minihalo and output their positions and other physical quantities at $z=49$. To enhance the numerical resolution, we perform particle splitting on DM and gas particles within a region that sufficiently encompasses their spatial distribution, roughly twice the final virial radius. We adopt the particle splitting technique described in [31], which arranges the $13$ child particles on a hexagonal close-packed array. We first run model L1 from the initial conditions generated at $z=49$ in the above-described manner, until the target gas cloud collapses ($\rho_{\rm b}=10^{6}\,{\rm GeV\,cm^{-3}}$). We then apply the same procedure to generate the initial conditions for the next level. We repeat this three times (from L0 to L3; see table 1). The highest-resolution model (L3) achieves a mass resolution of DM component $m_{\rm dm,min}=3.66\times 10^{-4}\,M_{\odot}$ with gravitational softening of $\varepsilon_{\rm min}=0.022\,$pc.
@@ 3 Numerical simulations
Figure 2: Density distributions of DM ($\rho_{\rm dm}$; top panels) and baryonic ($\rho_{\rm b}$; bottom panels) components projected on a cross-section through the density center. The left, center, and right panels show plots over $1000$, $10$, and $1$ pc per side, respectively.
@@ 3.1 Radial density profile
Figure 4: Radial profiles of the enclosed masses of baryon and DM components for Model L3 when $\rho_{\rm b,max}=10^{3}$ and $10^{9}\,{\rm GeV\,cm^{-3}}$. The dashed line shows a correlation as $M_{\rm enc}\propto r^{1.5}$.
@@ 3.3 DM core at $r<R_{\rm core}$
Before the onset of gravitational run-away collapse of the primordial gas cloud, a flat DM “core” with $\sim\!1$ pc develops. As shall be discussed later in section 3.4, the DM core consists of low velocity and low angular momentum materials which, however, cannot contract infinitesimally without any dissipation process for DM. Figure 5 plots the rotational-to-Keplerian velocity ratios calculated from the total mass, $v_{\rm rot}/v_{\rm Kepler,dm+b}$ (red), and from the DM mass alone, $v_{\rm rot}/v_{\rm Kepler,dm}$ (blue). At the early stage ($\rho_{\rm b,max}=10^{3}\,{\rm GeV\,cm^{-3}}$; dashed) both ratios exceed unity for $r<R_{\rm core}$. The near Keplerian rotation prevents further DM infall and sets the cored profile. [^3] This mechanism is different in origin from the core formation induced by baryonic feedback discussed in the context of galaxy formation. As baryons continue to contract (by dissipation), their cumulative mass inside $R_{\rm core}$ grows ($\rho_{\rm b,max}=10^{9}\,{\rm GeV\,cm^{-3}}$; solid). Consequently, $v_{\rm rot,dm}/v_{\rm Kepler,dm+b}$ falls to $\approx 1$, the DM particles experience adiabatic contraction, and the inner slope steepens once more [19]. Since $v_{\rm rot,dm}/v_{\rm Kepler,dm}$ remains $>1$, it is evident that the DM collapse is induced solely by the baryon’s gravitational potential.
Figure 6: Radial profiles of the specific angular momentum Panel (a): the specific angular momentum of baryon and DM components for Model L3 when $\rho_{\rm b}=10^{3}$ and $10^{9}\,{\rm GeV\,cm^{-3}}$. The gray lines show the specific angular momentum in the case of Keplerian rotation as $v_{\rm Kepler,dm+b}=\sqrt{G(M_{\rm enc,dm}(r)+M_{\rm enc,b}(r))/r}$. Panel (b): time evolution of the specific angular momentum of the selected three DM ($N$-body) particles calculated around the maximum density coordinates at each time. The filled circles represent the final states for three particles at the end of the simulation. The gray solid and thin lines show the radial profile of the DM component (the blue lines in panel a).
@@ 3.5 Fitting functions
$\quad\text{with}\quad\begin{cases}\{\rho_{i}\}=\{1.2\times 10^{5},1.2\times 10^{4},3.5\times 10^{3},3.5\times 10^{3}\}\,{\rm GeV\,cm^{-3}}\,,\\ \{\gamma_{i}\}=\{2.3,1.75,1.2,0.6\}\,,\end{cases}$ (3.1)
@@
We overlay the best-fit models $\rho_{\rm dm,fit}$ and $\rho_{\rm b,fit}$ (thin lines) for comparison in figure 3. Panel (c) in figure 3 shows the relative deviation between the fitting functions and the simulation results. The relative deviation in the DM density is confined within ±$\pm 10\%$. Although the relative deviation for the baryon density is somewhat larger (within ±$\pm 30\%$), this is likely due to the presence of substructure (see right bottom panel of figure 2).
@@ 5 Conclusion
    Is the resulting profile robust to resolution?: Multi-resolution tests (L0-L3) show that the dark-matter profile stabilizes long before numerical limits are reached: successive 13 times increases in particle number leave the core radius, inner slope, and overall multi-power-law shape essentially unchanged, and even the shallowest segment (0.01absent0.01$\approx$ 0.01≈ 0.010.10.10.10.1 pc) differs by only a few percent between runs. A four-segment fit reproduces the L3 profile to ±$\pm 10\%$ and the baryonic profile to ±$\pm 30\%$, the latter scatter arising from small-scale substructure rather than resolution errors. Because the central density and slope asymptote to the same values as resolution improves, the flattening and subsequent steepening of the cusp seen in L3 are physical, not numerical, confirming that our key conclusions are robust against variations in particle mass and softening length.
@@ Appendix A Numerical convergence
The density profiles obtained with our highest resolution model (L3) agree with those from lower-resolution ones, except very close to the resolution limit. By resolving the DM minihalo with progressively higher particle counts, the inferred central density and slope asymptote to stable values. The change of the DM density slope in model L3 is therefore a physical result, not a numerical artifact. Indeed, the shallow density structure that develops (of order 0.01similar-toabsent0.01$\sim$ 0.01∼ 0.010.10.10.10.1 pc in radius) persists with only minor differences when we vary the mass resolution or softening length.
@@ Acknowledgments
Numerical computations were performed on Cray XC50 and XD2000 at CfCA in the National Astronomical Observatory of Japan and Yukawa-21 at YITP in Kyoto University. Numerical analyses were carried out on analysis servers at Center for Computational Astrophysics, National Astronomical Observatory of Japan. This work made use of gnuplot, numpy [40] , scipy [41], and matplotlib [42] software packages. This work was supported by JSPS KAKENHI Grant Numbers JP21H01123 and JP23K20864 (S.H.).

### v2

@@ 1 Introduction
Gravitational wave (GW) astronomy offers a new avenue to probe fundamental physics [4]. Dynamical friction exerted by ambient DM on a black‑hole (BH) binary perturbs its orbital evolution and imprints a characteristic phase shift on the emitted GW signal [5]. If the central BH is surrounded by concentrated DM, dubbed a spike [6], dress [7], or mound [8, 9], the accumulated phase shift can be significant. State‑of‑the‑art waveform calculations suggest that next‑generation facilities are expected to have enough sensitivities to detect the effects [10, 11, 12]; e.g., third-generation ground-based (ET/CE) and space-based (LISA-like) detectors. Future GW observatories can thus serve as indirect DM detectors, but the feasibility heavily relies on the model of the DM density profile.
@@
The first generation of stars (Population III stars; hereafter Pop III stars) are thought to be formed at the centers of DM minihalos with masses $\sim\!10^{5}$$10^{6}\,M_{\odot}$ at redshifts $z\sim 30$$15$ [18, 19]. Massive Pop III stars collapse directly into BHs, which can become, under some suitable conditions, as massive as $10^{3}$$10^{5}\,M_{\odot}$. Because both the baryons and the nascent BH reside deep in the potential well, a dense DM cusp is expected to develop through adiabatic contraction [20, 21]. Such DM overdensities may substantially affect the GW signal associated with the merger of Pop III remnant BHs.
We present the result of very high resolution cosmological hydrodynamical simulations that resolve the DM distribution inside a minihalo without performing particle splitting from linear density perturbations $z=49$ to gravitational collapse. Our flagship run (labeled L3) attains a DM particle mass of $m_{\rm dm,min}=3.66\times 10^{-4}\,M_{\odot}$, and thus improves mass resolution by a factor $\approx 130$ relative to the previous works [26, 27]. We find that initially a slowly rotating core is formed. As the baryonic cloud contracts and the local Keplerian velocity increases, the DM core undergoes further contraction, developing a steeper inner slope. We derive a semi‑analytic formula for the DM density distribution at the Pop III star formation by fitting the simulation result and by extrapolating to protostellar sizes. The result provides a realistic input for GW waveform models with the contribution of concentrated DM.
2 Methods
@@
[…]
We adopt a two percent of the initial inter-particle distance as the gravitational softening length as $\varepsilon=L_{\rm box}/N_{\rm eff}^{1/3}/50$. We run model L0 until the baryon density reaches $\rho_{\rm b}=10^{6}\,{\rm GeV\,cm^{-3}}$, when the primordial gas cloud forms and collapses. At this stage, the DM minihalo hosting the gas cloud has a virial radius of approximately $R_{\rm virial}\simeq 100\,$pc and a mass of $M_{\rm virial}\simeq 3\times 10^{5}\,M_{\odot}$ at redshift $z_{\rm col}\simeq 22$, representing the typical properties of a minihalo for the Pop III star formation [18, 32].
@@
Table 1: Column 1: model name; Column 2: mass of the finest DM (NN-body) particle; Column 3: mass of the finest baryon/gas (SPH) particle; Column 4: gravitational softening length of the finest DM particle; and Column 5: redshift when the gas cloud collapses ($\rho_{\rm b,max}=10^{6}\,{\rm GeV\,cm^{-3}}$).
Next, we mark the DM and gas particles within the target minihalo and output their positions and other physical quantities at $z=49$. To enhance the numerical resolution, we perform particle splitting on DM and gas particles within a region that sufficiently encompasses their spatial distribution, roughly twice the final virial radius. We adopt the particle splitting technique described in [33], which arranges the $13$ child particles on a hexagonal close-packed array. Using the same L0 snapshot at $z=49$ as a common progenitor, we build three independent higher-resolution initial-condition sets by splitting each DM and gas particle into $13^{n}$ child particles ($n=1,2,3$). We then evolve these three sets separately as models L1–L3 (table 1) until the target gas cloud collapses ($\rho_{\rm b}=10^{6}\,{\rm GeV\,cm^{-3}}$). The highest-resolution model (L3) achieves a mass resolution of DM component $m_{\rm dm,min}=3.66\times 10^{-4}\,M_{\odot}$ with gravitational softening of $\varepsilon_{\rm min}=0.022\,$pc.
@@ 3 Numerical simulations
Figure 2: Density distributions of DM ($\rho_{\rm dm}$; top panels) and baryonic ($\rho_{\rm b}$; bottom panels) components projected on a cross-section through the density center. The left, center, and right panels show plots over $1000$, $10$, and 11 pc per side, respectively.
@@ 3.1 Radial density profile
Figure 4: Radial profiles of the enclosed masses of baryon and DM components (red and blue lines) for Model L3 when $\rho_{\rm b,max}=10^{3}$ and $10^{9}\,{\rm GeV\,cm^{-3}}$ (dashed and solid line styles). The long-dashed line shows a correlation as $M_{\rm enc}\propto r^{1.5}$.
@@ 3.3 DM core at $r<R_{\rm core}$
Before the onset of gravitational run-away collapse of the primordial gas cloud, a flat DM “core” with $\sim\!1$ pc develops. As shall be discussed later in section 3.4, the DM core consists of low velocity and low angular momentum materials which, however, cannot contract infinitesimally without any dissipation process for DM. Figure 5 plots the rotational-to-Keplerian velocity ratios calculated from the total mass, $v_{\rm rot}/v_{\rm Kepler,dm+b}$ (red), and from the DM mass alone, $v_{\rm rot}/v_{\rm Kepler,dm}$ (blue). At the early stage ($\rho_{\rm b,max}=10^{3}\,{\rm GeV\,cm^{-3}}$; dashed) both ratios exceed unity for $r<R_{\rm core}$. The near Keplerian rotation prevents further DM infall and sets the cored profile. [^3] This mechanism is different in origin from the core formation induced by baryonic feedback discussed in the context of galaxy formation. As baryons continue to contract (by dissipation), their cumulative mass inside $R_{\rm core}$ grows ($\rho_{\rm b,max}=10^{9}\,{\rm GeV\,cm^{-3}}$; solid). Consequently, $v_{\rm rot,dm}/v_{\rm Kepler,dm+b}$ falls to $\approx 1$, the DM particles experience adiabatic contraction, and the inner slope steepens once more [20]. Since $v_{\rm rot,dm}/v_{\rm Kepler,dm}$ remains >1>1, it is evident that the DM collapse is induced solely by the baryon’s gravitational potential.
In the late, high-density phase, the DM inside the previously formed cored profile undergoes a baryon-induced re-contraction: as the dissipative baryon collapse deepens the central potential, the DM responds adiabatically. Because DM lacks dissipation, this re-contraction is subordinate to the baryon collapse rather than self-driven. Consequently, the DM density center follows the baryon density center, and within our spatial resolution, the two peaks coincide during this phase (see figure 3). This confirms in our simulations that a DM core established at earlier times can steepen again once the baryon collapse proceeds.
Figure 6: Radial profiles of the specific angular momentum Panel (a): the specific angular momentum of baryon and DM components for Model L3 when $\rho_{\rm b}=10^{3}$ and $10^{9}\,{\rm GeV\,cm^{-3}}$. The gray lines show the specific angular momentum in the case of Keplerian rotation as $v_{\rm Kepler,dm+b}=\sqrt{G(M_{\rm enc,dm}(r)+M_{\rm enc,b}(r))/r}$. Panel (b): time evolution of the specific angular momentum of the selected three DM (NN-body) particles calculated around the maximum density coordinates at each time. The filled circles represent the final states for three particles at the end of the simulation. The gray solid and thin lines show the radial profile of the DM component (the blue lines in panel a).
@@ 3.5 Fitting functions
[…]
@@
We overlay the best-fit models $\rho_{\rm dm,fit}$ and $\rho_{\rm b,fit}$ (thin lines) for comparison in figure 3. Panel (c) in figure 3 shows the relative deviation between the fitting functions and the simulation results. The relative deviation in the DM density is confined within $\pm 10\%$. Although the relative deviation for the baryon density is somewhat larger (within $\pm 30\%$), this is likely due to the presence of substructure (see right bottom panel of figure 2).
@@ 5 Conclusion
    Is the resulting profile robust to resolution?: Multi-resolution tests (L0-L3) show that the dark-matter profile stabilizes long before numerical limits are reached: successive 13 times increases in particle number leave the core radius, inner slope, and overall multi-power-law shape essentially unchanged, and even the shallowest segment ($\approx 0.01$$0.1$ pc) differs by only a few percent between runs. A four-segment fit reproduces the L3 profile to $\pm 10\%$ and the baryonic profile to $\pm 30\%$, the latter scatter arising from small-scale substructure rather than resolution errors. Because the central density and slope asymptote to the same values as resolution improves, the flattening and subsequent steepening of the cusp seen in L3 are physical, not numerical, confirming that our key conclusions are robust against variations in particle mass and softening length.
@@ Appendix A Numerical convergence
The density profiles obtained with our highest resolution model (L3) agree with those from lower-resolution ones, except very close to the resolution limit. By resolving the DM minihalo with progressively higher particle counts, the inferred central density and slope asymptote to stable values. The change of the DM density slope in model L3 is therefore a physical result, not a numerical artifact. Indeed, the shallow density structure that develops (of order $\sim 0.01$$0.1$ pc in radius) persists with only minor differences when we vary the mass resolution or softening length.
@@ Acknowledgments
The authors thank Jonathan Tan and Jesus Zavala for fruitful discussion. Numerical computations were performed on Cray XC50 and XD2000 at CfCA in the National Astronomical Observatory of Japan and Yukawa-21 at YITP in Kyoto University. Numerical analyses were carried out on analysis servers at Center for Computational Astrophysics, National Astronomical Observatory of Japan. This work made use of gnuplot, numpy [42] , scipy [43], and matplotlib [44] software packages. This work was supported by JSPS KAKENHI Grant Numbers JP21H01123 and JP23K20864 (S.H.).